PLEASE HELP ME! I'M BEGGING YOU!
Solve the following the substitution. Write your answer in the form of an ordered pair.
1. x + 5y = - 8
x = 3y
2. 3x - 3y = 12
3x + y = 8
3. 4x + 3y = 7
x = 6y - 5
4. y = 3x + 3
4x + 3y = -17
5. a = 4b
5a + 4b = 12

Answer by jim_thompson5910(35256)

# 1
Hint:

x + 5y = - 8

3y + 5y = - 8 ... replace x with 3y

Solve for y. Then use this solution to find the solution for x.
-------------------------------------------------------
# 2
Solve the second equation 3x + y = 8 for y to get y = -3x+8

Then move back the first equation 3x - 3y = 12 and replace y with -3x+8 to get

3x - 3(-3x + 8) = 12

Solve that for x. Then use the solution to find y.

Let me know if this helps. Thanks.